# Nondeterminism-Aware Optimistic Verification for Floating-Point Neural Networks

### v2

@@ 2.2. Benign numerical drift is ubiquitous and regular
Figure 3. NAO overview. Phase 0 (Model Setup): model owner commits to weights, graph, empirical error percentile thresholds, and per-operator trace schema. Phase 1 (Optimistic Execution): proposer posts a commitment for $(x,y)$. Phase 2 (Dispute Localization): an NN-way, Merkle-anchored dispute game recursively narrows disagreement to a single operator following a canonical partition policy. Phase 3 (Single-Operator Adjudication): either a sound IEEE-754 error-bound check or a committee vote against empirical thresholds decides the outcome, triggering payment or slashing.:
2.3. Sound theory alone is too loose (and exploitable)
Given the small empirical errors above, a natural baseline is to instantiate a compositional framework for theoretical IEEE-754 rounding bounds. For each operator with its input, we compute a worst-case, element-wise rounding envelope that is sound under standard floating-point error models. However, theory-only envelopes are far looser than what we observe in practice (comparison in Fig. 6). This slack enlarges the attack surface. With a traced model and per-op theoretical budgets in hand, an adversary can inject perturbations within these element-wise bounds after each op. Using projected-gradient steps on the task loss, they can steer the cumulative budget along sensitive directions to induce label flips or semantic drift (Fig. 1). On large models, such “theory-only” checks admit a small but non-trivial attack success rate (Table 2), demonstrating that worst-case soundness alone is insufficient for economically robust verification.
Bridge: closing the gap. The real problem therefore demands new ideas: eliminating the exploitable slack those bounds leave in real deployments. NAO closes this gap by calibrating empirical, per-model, per-operator error percentile thresholds across heterogeneous hardware and using them to guide dispute localization; leaves are then adjudicated cheaply via either certified theoretical checks or small committee votes against the stored empirical thresholds. This dual design embraces nondeterminism, removes replication burden, and makes verification practical and defensible. Later sections detail the mechanism.
@@ 2.3. Sound theory alone is too loose (and exploitable)
3. System architecture and protocol overview
System at a glance. NAO provides verifiable results for NN workloads by combining optimistic-rollup-style economics with tolerance-aware verification. Fig. 3 shows the four phases: in Phase 0 (Setup), the model owner commits Merkle roots of the weights and graph, and publishes calibrated per-operator empirical error percentile thresholds. In Phase 1 (Optimistic execution), a proposer runs the model to fulfill user’s request and the result finalizes after a short challenge window. If challenged, it runs a Merkle-anchored, threshold-guided interactive Phase 2 (dispute game), that recursively narrows disagreement from the entire model graph to a single PyTorch operator. Finally, Phase 3 (Single-operator adjudication) executes either (i) a sound, cheap but potentially permissive IEEE-754 theoretical-bound check, or (ii) a tighter but costlier committee vote against the empirical error percentile thresholds. Violations result in slashing.
@@ 3.1. Roles and assumptions
$\bullet$ On-chain contract: escrows payments, records commitments, enforces timeouts, and orchestrates the state machine.
@@
Assumptions. We assume IEEE-754 compliance with round-to-nearest-even; fast-math is disabled; operator libraries may reorder reductions but must respect IEEE semantics. When invoked, the committee has an honest majority, and no trusted hardware is required.
@@ 3.2. End-to-end protocol lifecycle
Phase 0: Model setup. At deployment time, the model owner serializes the PyTorch model into an acyclic dataflow graph $G=(V,E)$ with a canonical topological order, where each node $v_{i}\in V$ denotes a tensor operator (e.g., MatMul, Softmax, GELU) and each edge $(u,v)\in E$ captures a data dependency. The model state is then committed on-chain by merkleizing the weight tensors to obtain root $r_{w}$ and, separately, merkleizing the graph topology (operator signatures) to obtain root $r_{g}$. The pair $(r_{w},r_{g})$ is recorded on the blockchain, enabling lightweight provenance proofs for any subgraph that arises during later dispute phases (Sec. 6.2). We also perform an offline calibration (Sec. 4.2) to quantify cross-hardware floating-point variability per operator, yielding empirical error percentile thresholds, which are also committed as a root $r_{e}$. These commitments enable verifiable subgraph extraction: any subgraph $S\subseteq G$ carries membership proofs into $r_{g}$ and $r_{w}$.
@@
Phase 2: Dispute localization (interactive decomposition). If any verifier recomputes $y^{\prime}=G(x)$ and finds that the discrepancy between $y$ and $y^{\prime}$ exceeds the empirical error percentile threshold, they may initiate a dispute game and act as a challenger. The contract then freezes proposer’s collateral $\sigma_{P}$, starts a dispute game, and initializes round k=0k{=}0 with $S_{0}{:=}G$. Each round proceeds (details in Sec. 6.3):
@@
$\bullet$ Advance. The contract updates the dispute game state: sets $S_{k+1}\leftarrow S_{k}^{j^{\star}}$ and increments $k$, and enforces per-round timeouts. Failure to act within the timeout loses that round.
@@
Phase 3: Single-operator adjudication. At the leaf, both parties agree on the operator type $v^{*}$ and the inputs $a$. The challenger then tests for fraud using one of two methods: (i) a theoretical-bound check, where a verifiable VM computes element-wise FP error bounds (Sec. 4.1): If the proposer’s claimed outputs fall outside these bounds for any element, the contract slashes $\sigma_{P}$ and rewards the challenger. (ii) When the bound is inconclusive (the claim lies within the theoretical error bound), a committee check, where multiple independent runners execute $(v^{*},a)$ and vote using the pre-calibrated empirical error percentile thresholds: this test is much tighter but more costly.
@@ 4.1. Computing theoretical error bound
We adopt the standard floating-point model as follows: for ${+,,×,/}\circ\in\{$$\times$,/\}, $\operatorname{fl}(x\circ y)=(x\circ y)(1+\delta)$ with $|\delta|\leq u$ ($u$ is unit roundoff). For library intrinsics we use vendor-stated maximum-ULP errors for CUDA math functions (Guide, 2013). For reductions composed of $k$ basic ops, we support both worst-case deterministic  (Higham, 2002; Goldberg, 1991) and relatively tighter probabilistic (Madry et al., 2018; Higham and Mary, 2019) bounds. A canonical example is sequential summation $s=\sum_{i=1}^{n}x_{i}$ (k=n1k=n-1):
@@
FX-based co-execution and operator templates. We extend PyTorch FX (PyTorch Contributors, 2025) interpreter to co-execute values and bounds. Each traced operator is lowered to a short sequence of primitives (adds/multiplications/reductions/intrinsics). For a primitive $f(x_{1},\dots,x_{m})$ with current inputs $\{x_{i}\}$ and already-accumulated intra-operator input errors $\{\varepsilon_{x_{i}}\}$, we use a first-order sensitivity envelope
@@
Figure 4. Comparison between deterministic and probabilistic theoretical error bounds for different operator types in Qwen-8B and BERT-large. :
Putting it together. Our library of operator templates yields $\tau_{v}^{\text{theo}}$ cheaply during a single traced run. These per-operator bounds are later used for single-operator verification. Fig. 4 contrasts deterministic $\gamma_{k}$ with probabilistic $\widetilde{\gamma}_{k}$ across representative operators; the probabilistic bounds are markedly tighter, especially at large $k$. However, worst-case $\gamma_{k}$ and even $\widetilde{\gamma}_{k}$ remain conservative for real ML workloads where mixed signs induce cancellation; empirically, observed cross-device discrepancies are often orders of magnitude smaller. We therefore pre-calibrate per-operator error percentile profiles on representative inputs to construct empirical thresholds and use them alongside the theoretical bounds.
@@ 4.2. Calibrating empirical error percentile threshold
The set $\{\tau_{\mathrm{abs}}^{(i)},\tau_{\mathrm{rel}}^{(i)}\}_{i=1}^{|V|}$ is bundled with the model commitment $(r_{w},r_{g})$ and is included in the on-chain commitment.
@@ 5. Threat model and attack design
This section formalizes the adversary and our security objectives, and states that in NAO the only attack surface is the two error bounds. We then describe the attacks used to probe these interfaces under their respective conditions and, finally, demonstrate the robustness of the error bounds.
@@ 5.1. Adversary and objectives
Adversary capabilities. We consider only the proposer adversarial; all other parties follow the protocol. The adversary is white-box: they know model commitment, input, canonical topological order, empirical error percentile thresholds, and the dispute-partition policy. The adversary may alter intermediate tensors during the computation but cannot change weights or graph topology without violating Merkle commitments, nor influence committee sampling or cryptography.
@@ 5.2. Attack surface and empirical headroom
The attack surface is numerical: at selected $v$, the proposer injects additive perturbations $h_{v}\leftarrow h_{v}+\Delta_{v}$ to change the final decision while remaining within challenger’s admissible sets. To quantify headroom, we trace mean cross-device error per operator and plot it against normalized depth (Fig. 2). Across BERT-large, Qwen-8B, and ResNet-152, errors are essentially flat with localized spikes: we observe no systematic accumulation of error as depth increases. Typical magnitudes are $10^{-6}$-$10^{-5}$. Thus, flipping outputs while staying admissible requires coordinating many sub-micro deviations; over-allocation will be isolated by the dispute game.
@@ 5.4. Optimization-based attack method
We implement projected-gradient attack (Madry et al., 2018) over $\{\Delta_{v}\}$ to flip the predicted class/token while staying within a chosen admissible set.
(10) $L_{\text{margin}}(z)\;=\;z_{c_{2}}-z_{c_{1}},$
(12) $\Delta_{v}^{t+1}=\mathrm{reshape}\;\!\big(s\odot a^{\star}\big).$
@@ 5.5. Attack evaluation
Datasets and hyperparameters. For BERT-large (Devlin et al., 2019) we use DBpedia (Auer et al., 2007); for Qwen3-8B (Team, 2025) we use C4 (Raffel et al., 2019), feeding the first 10% of each sequence and targeting the next token; for ResNet-152 (He et al., 2016) we use ImageNet (Deng et al., 2009) validation set. PGD/Adam and projection operators follow Sec. 5.4. Early stopping triggers when the last 10 updates satisfy $|m_{t}-m_{t-1}|<10^{-3}|m_{0}|$ and $|m_{t}|<10^{-3}|m_{0}|$ ($m_{t}$ is the logit margin at time step $t$). We use 500 samples (2500 attack targets) for the theoretical-bound attack and 250 (1250 attack targets) for the empirical-threshold attack to balance statistical precision and computational cost.
@@
Figure 5. Normalized margin change on failed attacks ($\alpha{=}1$). Empirical thresholds yield near-zero progress across models; theoretical bounds(p) show heavier tails for the LLM. Numbers above boxes show sample sizes.:
@@
6. NAO: Optimistic operator-level dispute protocol
In this section, we adopt an optimistic design in which a public L1 blockchain serves as arbiter. A proposer commits to the model, inputs, and outputs; if no challenger disputes within a window, the result finalizes cheaply on L1. Upon dispute, an interactive localization game narrows disagreement to a single operator guided by the error tolerance, after which L1 performs a constant-cost check; the losing party is slashed.
@@ 6. NAO: Optimistic operator-level dispute protocol
Prior verification systems, including refereed delegation of computation, which guarantees correctness by querying multiple untrusted servers and resolving disagreements via a referee (Canetti et al., 2013, 2011), and interactive proofs (Goldwasser et al., 2008; Kalai et al., 2022). Rollup-style verification games (e.g., Truebit, Arbitrum, OP Stack (Teutsch and Reitwießner, 2024; Kalodner et al., 2018; Optimism, 2025b, a)) instantiate this over CPU-style VMs via instruction-level bisection. NN on heterogeneous accelerators violates both: GPU kernels cannot be stepped by existing VMs and floating-point non-associativity plus parallel scheduling introduce cross-hardware variability (PyTorch Contributors, 2024; Higham, 2002).
NAO preserves the optimistic, interactive structure but replaces the units of dispute and the base-case check: (i) we perform N-way partition on a Merkle-anchored, operator-granular NN graph rather than a VM trace; (ii) we guide partition selection via tolerance-aware error thresholds calibrated per operator; and (iii) at the leaf we perform either a sound IEEE-754 theoretical-bound check or (when bounds are permissive) a committee vote against empirical error percentile thresholds. This design is new in (a) moving from instruction-level to operator-level partition, (b) defining a deterministic selection rule under non-deterministic floats, and (c) providing a constant-cost L1 adjudication path that respects IEEE semantics without requiring a GPU VM.
@@ 6.2. Operator-granular graph representation and verifiable subgraph extraction
Merkle commitments to weights, graph, and interfaces. We commit separately to weights and graph structure via Merkle trees whose roots are published on chain. We define $\mathsf{H}(\cdot)$ as the SHA-256 hash function.
@@
captures kind, attributes, and topology (edges are implied by arguments). The leaf hash is $\mathcal{H}_{g}(n)=\mathsf{H}\big(\sigma(n)\big)$, forming the structure tree $\mathcal{T}_{g}$ with root $r_{g}$. This construction enables logarithmic-depth proofs, making on-chain verification feasible even for large models.

### v3

@@ 2. System architecture and protocol overview
Figure 2. TAO overview. Phase 0 (Model Setup): model owner commits to weights, graph, empirical error percentile thresholds, and per-operator trace schema. Phase 1 (Optimistic Execution): proposer posts a commitment for $(x,y)$. Phase 2 (Dispute Localization): an NN-way, Merkle-anchored dispute game recursively narrows disagreement to a single operator following a canonical partition policy. Phase 3 (Single-Operator Adjudication): either a sound IEEE-754 error-bound check or a committee vote against empirical thresholds decides the outcome, triggering payment or slashing.
A diagram showing the workflow of the system. It consists of four phases: Model Setup, Optimistic Execution, Dispute Localization, and Single-Operator Adjudication. Each phase is described in detail, highlighting the key steps and interactions between the proposer, challenger, committee members, and users.
@@
System at a glance. TAO provides verifiable results for NN workloads by combining optimistic dispute-based verification with tolerance-aware verification. Fig. 2 shows the four phases: in Phase 0 (Setup), the model owner commits Merkle roots of the weights and graph, and publishes calibrated per-operator empirical error percentile thresholds. In Phase 1 (Optimistic execution), a proposer runs the model to fulfill user’s request and the result finalizes after a short challenge window. If challenged, it runs a Merkle-anchored, threshold-guided interactive Phase 2 (dispute game), that recursively narrows disagreement from the entire model graph to a single PyTorch operator. Finally, Phase 3 (Single-operator adjudication) executes either (i) a sound, cheap but potentially permissive IEEE-754 theoretical-bound check, or (ii) a tighter but costlier committee vote against the empirical error percentile thresholds. Violations result in slashing.
@@ 2.1. Roles and assumptions
$\bullet$ Coordinator: an authenticated coordination service that records commitments, enforces timeouts, and orchestrates the state machine.
@@
Assumptions. We assume IEEE-754 compliance with round-to-nearest-even; operator libraries may reorder reductions but must respect IEEE semantics. When invoked, the committee has an honest majority, and no trusted hardware is required. Most specific optimizations are compatible with TAO, with details discussed in Sec. 7. We assume challengers have access to comparable compute resources as proposers, while the committee can have substantially less compute. Challengers are not assumed to be honest.
@@ 2.2. End-to-end protocol lifecycle
Phase 0: Model setup. At deployment time, the model owner serializes the PyTorch model into an acyclic dataflow graph $G=(V,E)$ with a canonical topological order, where each node $v_{i}\in V$ denotes a tensor operator (e.g., MatMul, Softmax, GELU) and each edge $(u,v)\in E$ captures a data dependency. The model state is then committed by merkleizing the weight tensors to obtain root $r_{w}$ and, separately, merkleizing the graph topology (operator signatures) to obtain root $r_{g}$. The pair $(r_{w},r_{g})$ is recorded on the coordinator backend, enabling lightweight provenance proofs for any subgraph that arises during later dispute phases (Sec. 5.2). We also perform an offline calibration (Sec. 3.2) to quantify cross-hardware floating-point variability per operator, yielding empirical error percentile thresholds, which are also committed as a root $r_{e}$. These commitments enable verifiable subgraph extraction: any subgraph $S\subseteq G$ carries membership proofs into $r_{g}$ and $r_{w}$.
@@
Phase 2: Dispute localization (interactive decomposition). If any verifier recomputes $y^{\prime}=G(x)$ and finds that the discrepancy between $y$ and $y^{\prime}$ exceeds the empirical error percentile threshold, they may initiate a dispute game and act as a challenger. The coordinator then freezes proposer’s collateral $\sigma_{P}$, starts a dispute game, and initializes round k=0k{=}0 with $S_{0}{:=}G$. Each round proceeds (details in Sec. 5.3):
@@
$\bullet$ Advance. The coordinator backend updates the dispute game state: sets $S_{k+1}\leftarrow S_{k}^{j^{\star}}$ and increments $k$, and enforces per-round timeouts. Failure to act within the timeout loses that round.
@@
Phase 3: Single-operator adjudication. At the leaf, both parties agree on the operator type $v^{*}$ and the inputs $a$. The challenger then tests for fraud using one of two methods: (i) a theoretical-bound check, where a verifiable VM computes element-wise FP error bounds (Sec. 3.1): If the proposer’s claimed outputs fall outside these bounds for any element, the coordinator confiscates $\sigma_{P}$ and rewards the challenger. (ii) When the bound is inconclusive (the claim lies within the theoretical error bound), a committee check, where multiple independent runners execute $(v^{*},a)$ and vote using the pre-calibrated empirical error percentile thresholds: this test is much tighter but more costly.
@@ 3.1. Computing theoretical error bound
We adopt the standard floating-point model as follows: for ${+,,×,/}\circ\in\{$$\times$,/\}, $\operatorname{fl}(x\circ y)=(x\circ y)(1+\delta)$ with $|\delta|\leq u$ ($u$ is unit roundoff). For library intrinsics we use vendor-stated maximum-ULP errors for CUDA math functions (guide2013cuda). For reductions composed of $k$ basic ops, we support both worst-case deterministic  (higham2002accuracy; goldberg1991every) and relatively tighter probabilistic (madry2018towards; higham2019new) bounds. A canonical example is sequential summation $s=\sum_{i=1}^{n}x_{i}$ (k=n1k=n-1):
@@
FX-based co-execution and operator templates. We extend PyTorch FX (pytorch-fx) interpreter to co-execute values and bounds. Each traced operator is lowered to a short sequence of primitives (adds/multiplications/reductions/intrinsics). For a primitive $f(x_{1},\dots,x_{m})$ with current inputs $\{x_{i}\}$ and already-accumulated intra-operator input errors $\{\varepsilon_{x_{i}}\}$, we use a first-order sensitivity envelope
@@
Figure 3. Comparison between deterministic and probabilistic theoretical error bounds for different operator types in Qwen-8B and BERT-large.
A graph comparing deterministic and probabilistic theoretical error bounds for different operator types in Qwen-8B and BERT-large. The x-axis represents the node types, while the y-axis the mean of absolute error. The graph shows that probabilistic bounds are significantly tighter than deterministic bounds.
Putting it together. Our library of operator templates yields $\tau_{v}^{\text{theo}}$ cheaply during a single traced run. These per-operator bounds are later used for single-operator verification. Fig. 3 contrasts deterministic $\gamma_{k}$ with probabilistic $\widetilde{\gamma}_{k}$ across representative operators; the probabilistic bounds are markedly tighter, especially at large $k$. These bounds provide a cheap, hardware-agnostic “ground-truth” leaf check that doesn’t rely on committee honesty. Many real-world misbehaviors (e.g., model swaps, undeclared quantization) typically deviate far outside theoretical bounds and can be rejected quickly.
However, worst-case $\gamma_{k}$ and even $\widetilde{\gamma}_{k}$ remain conservative for real ML workloads where mixed signs induce cancellation; empirically, observed cross-device discrepancies are often orders of magnitude smaller. We therefore pre-calibrate per-operator error percentile profiles on representative inputs to construct empirical thresholds and use them alongside the theoretical bounds.
@@ 3.2. Calibrating empirical error percentile threshold
The set $\{\tau_{\mathrm{abs}}^{(i)},\tau_{\mathrm{rel}}^{(i)}\}_{i=1}^{|V|}$ is bundled with the model commitment $(r_{w},r_{g})$ and is included in the model commitment recorded by the coordinator.
@@ 4. Threat model and attack design
This section evaluates adversarial proposer behavior under the protocol’s numerical admissible sets; it does not attempt to model all deployment risks (e.g., coordinator compromise). We then describe one of the attack methods used to probe these interfaces under their respective conditions and, finally, demonstrate the robustness of the error bounds.
@@ 4.1. Adversary and objectives
Adversary capabilities. In this section, we consider the case where only the proposer is adversarial; all other parties follow the protocol. The adversary is white-box: they know model commitment, input, canonical topological order, empirical error percentile thresholds, and the dispute-partition policy. The adversary may alter intermediate tensors during the computation but cannot change weights or graph topology without violating Merkle commitments, nor influence committee sampling or cryptography.
@@ 4.2. Attack surface and empirical headroom
The attack surface is numerical: at selected $v$, the proposer injects additive perturbations $h_{v}\leftarrow h_{v}+\Delta_{v}$ to change the final decision while remaining within challenger’s admissible sets. To quantify headroom, we trace mean cross-device error per operator and plot it against normalized depth (Fig. 4). Across BERT-large, Qwen-8B, and ResNet-152, errors are essentially flat with localized spikes: we observe no systematic accumulation of error as depth increases. Typical magnitudes are $10^{-6}$-$10^{-5}$. Thus, flipping outputs while staying admissible requires coordinating many extremely small deviations; over-allocation will be isolated by the dispute game.
@@ 4.4. Optimization-based attack method
We implement projected-gradient attack (madry2018towards) over $\{\Delta_{v}\}$ to flip the predicted class/token while staying within a chosen admissible set.
@@
(10) $L_{\text{margin}}(z)\;=\;z_{c_{2}}-z_{c_{1}},$
(12) $\Delta_{v}^{t+1}=\mathrm{reshape}\;\!\big(s\odot a^{\star}\big).$
@@ 4.5. Attack evaluation
Datasets and hyperparameters. For BERT-large (devlin2019bert) we use DBpedia (auer2007dbpedia); for Qwen3-8B (qwen3technicalreport) we use C4 (2019t5), feeding the first 10% of each sequence and targeting the next token; for ResNet-152 (he2016deep) we use ImageNet (deng2009imagenet) validation set. PGD/Adam and projection operators follow Sec. 4.4. Early stopping triggers when the last 10 updates satisfy $|m_{t}-m_{t-1}|<10^{-3}|m_{0}|$ and $|m_{t}|<10^{-3}|m_{0}|$ ($m_{t}$ is the logit margin at time step $t$). We use 500 samples (2500 attack targets) for the theoretical-bound attack and 250 (1250 attack targets) for the empirical-threshold attack to balance statistical precision and computational cost.
@@
Figure 5. Normalized margin change on failed attacks ($\alpha{=}1$). Empirical thresholds yield near-zero progress across models; theoretical bounds(p) show heavier tails for the LLM. Numbers above boxes show sample sizes.
A box plot showing the normalized margin change on failed attacks with alpha equal to 1. Empirical thresholds yield near-zero progress across models, while theoretical bounds show heavier tails for the LLM. Numbers above boxes indicate sample sizes.
@@
5. TAO: Optimistic operator-level dispute protocol
In this section, we adopt an optimistic design in which a coordinator acts as the adjudicator. A proposer commits to the model, inputs, and outputs; if no challenger disputes within a window, the result finalizes cheaply. Upon dispute, an interactive localization game narrows disagreement to a single operator guided by the error tolerance, after which the coordinator performs a constant-cost check; the losing party is penalized.
@@ 5. TAO: Optimistic operator-level dispute protocol
Prior verification systems, including refereed delegation of computation, which guarantees correctness by querying multiple untrusted servers and resolving disagreements via a referee (CanettiRivaRothblumIC13; CRR-CCS11), and interactive proofs (GKR08; KRR22). Rollup-style verification games (e.g., Truebit, Arbitrum, OP Stack (teutsch2024scalable; kalodner2018arbitrum; optimismcannonfpvm; optimismfaultproof)) instantiate this over CPU-style VMs via instruction-level bisection. NN on heterogeneous accelerators violates both: GPU kernels cannot be stepped by existing VMs and floating-point non-associativity plus parallel scheduling introduce cross-hardware variability (pytorch-reproducibility; higham2002accuracy).
TAO preserves the optimistic, interactive structure but replaces the units of dispute and the base-case check: (i) we perform N-way partition on a Merkle-anchored, operator-granular NN graph rather than a VM trace; (ii) we guide partition selection via tolerance-aware error thresholds calibrated per operator; and (iii) at the leaf we perform either a sound IEEE-754 theoretical-bound check or (when bounds are permissive) a committee vote against empirical error percentile thresholds. This design is new in (a) moving from instruction-level to operator-level partition, (b) defining a deterministic selection rule under non-deterministic floats, and (c) providing a constant-cost L1 adjudication path that respects IEEE semantics without requiring a GPU VM.
@@ 5.2. Operator-granular graph representation and verifiable subgraph extraction
Merkle commitments to weights, graph, and interfaces. We commit separately to weights and graph structure via Merkle trees whose roots are recorded by the coordinator. We define $\mathsf{H}(\cdot)$ as the SHA-256 hash function.
@@
captures kind, attributes, and topology (edges are implied by arguments). The leaf hash is $\mathcal{H}_{g}(n)=\mathsf{H}\big(\sigma(n)\big)$, forming the structure tree $\mathcal{T}_{g}$ with root $r_{g}$. This construction enables logarithmic-depth proofs, making fast verification feasible even for large models.